OpenStudy (beccab003):

Help with shape measurements! Medal to best answer. Thanks for the help!! Question will be posted below.

OpenStudy (beccab003):

If RZX = RST, calculate the perimeter of RST.

OpenStudy (beccab003):

I just don't understand how I am supposed to find the measurements of the other line lengths because I tried and the numbers came out weird.

OpenStudy (beccab003):

@Directrix @jim_thompson5910 @DanJS

jimthompson5910 (jim_thompson5910):

do you see what the length of YX and WT must be?

OpenStudy (beccab003):

yeah, 9 and 15. I get that part.

OpenStudy (danjs):

\[\frac{ 21 }{ 21+ZS } = \frac{ 9 }{ 15 }\] is that legal to do

jimthompson5910 (jim_thompson5910):

that's valid DanJs

OpenStudy (beccab003):

I tried to use that kind of method to solve this but without the left side part that you wrote.

OpenStudy (danjs):

if you solve that ZS=14 then use that in a new ratio

OpenStudy (danjs):

\[\frac{ 21 }{ 21+14 } = \frac{ 30 }{ 30+XT }\]

OpenStudy (danjs):

then you have all the parts to add up

OpenStudy (beccab003):

sweet! I know how to do the problem from there!!!

OpenStudy (beccab003):

it's 115!!! I just didn't know how to get that first number...THANKS!!

OpenStudy (danjs):

welcome